OpenStudy (anonymous):

mulitply then simplify by factoring if possible. Assume that all variables under radicals represent nonnegative numbers. Square root 11t square root 11t ill draw it

OpenStudy (anonymous):

\[\sqrt{11t} \sqrt{11t} =\]

OpenStudy (saifoo.khan):

\[\sqrt a * \sqrt a = a\]

OpenStudy (anonymous):

is that the answer

OpenStudy (saifoo.khan):

No. that's an example. The answer will be similar to that.

OpenStudy (anonymous):

so the answer would be \[\sqrt{11}*\sqrt{11} a = 121\]

OpenStudy (anonymous):

\[\sqrt{11t*11t}=11t\]

OpenStudy (saifoo.khan):

nope. Whenever you multiply 2 radicals which have same terms inside, the answer only be that's inside the radical. In other words the radicals cancels out.

OpenStudy (saifoo.khan):

Like @mustry did.

OpenStudy (anonymous):

okay so the answer is \[\sqrt{11t} *\sqrt{11t}= 11t\]

OpenStudy (anonymous):

actual ur r8

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

u@mizzkuntry

OpenStudy (saifoo.khan):

actual ur r8: Actually you're right. Lol.

OpenStudy (anonymous):

oh lol thank you

OpenStudy (anonymous):

i have two more questions can you help me wit them

OpenStudy (anonymous):

hahahhah okay